OpenStudy (anonymous):

. Raise the quantity in parentheses to the indicated exponent, and simplify the resulting expression. (-2x9^y^4)^4

OpenStudy (campbell_st):

use the index rule \[(x^a)^b = x^{a \times b}\] so your problem is \[(-2x^9 y^4)^4 = (-2)^4 \times x^{9 \times 4} \times y^{4\times4}\] I'll leave it to you to finish